# Scheduled Curiosity-Deep Dyna-Q: Efficient Exploration for Dialog Policy Learning

### v1

@@ 1 Introduction
Since human-computer interaction and natural language processing are in high demand in industry and daily life, and the task-oriented dialog system has become a hot topic and deserves further study and research. Dialog policy model is used to select the best action at each step of a dialog. In a task-oriented dialog system, the goal of the dialog is to complete a specific task, such as booking a movie. The dialog policy should select actions to efficiently achieve the goal. Dialog policy learning is often formulated as a reinforcement learning (RL) problem [1, 2], in which a dialog agent executes an action based on the observed state and receives a reward from the environment acted by real users. However, optimizing RL agents requires a large number of interactions with the environment, which is an expensive and time-consuming process. Therefore, how to grasp dialog policy efficiently within limited interactions remains a fundamental question for RL [3, 4, 5, 6, 7]. A common approach is to encourage the agent to explore the environment as sufficiently as possible through limited interactions. In contrast to supervised and unsupervised learning, reinforcement learning does not contain a large amount of training data with or without labels prepared in advance, but rather generates experience through the agent’s interactions with the environment and continually optimizes the agent based on feedback from the environment. Therefore, exploration of the environment is crucial for RL, and extensive research has been conducted on promoting agents to explore the environment. Count-based exploration is an intuitive approach in which each state observed by the agent is recorded and the agent is encouraged to learn states with fewer occurrences. The most evident drawback of this approach is its inability to handle complex and high-dimensional data. To address this, Bellemare et al. engineered a model that acquires the distribution of visited environment states to measure the novelty of states and encourage agents to observe unfamiliar states [8]. In addition, to deal with high-dimensional data, Tang et al. suggested mapping high-dimensional states to hash codes [9], and Liu et al. introduced embedding networks to encode the state space [10]. However, a common flaw in such approaches is that they do not lead agents to explore states that have never existed before. Another method is to model the intrinsic motivation of the agents. One ingenious strategy is to utilize curiosity to facilitate state exploration. Pathak et al. [11] proposed an Intrinsic Curiosity Module (ICM) to model the state prediction error, which depicts the uncertainty and improvement in RL as a curiosity reward to encourage unfamiliar dialog states. This research makes some simplifications and adjustments to ICM to make it more suitable for task-oriented dialog systems to guide unfamiliar dialog state exploration. This model leveraged the predictability of a state to depict its novelty. If the next state can be accurately predicted based on the current state and the upcoming action to be performed, it means that the action cannot lead the agent to see unfamiliar states, and it corresponds to a less curiosity reward. By contrast, if the next state is difficult to predict, the corresponding action receives a larger curiosity reward. The agent executes the action with the largest sum of curiosity rewards and rewards from the environment. Furthermore, the RL-based dialog agent faces another challenge. Generally, an episode of dialog based on RL is opened up by randomly selecting a user goal from the entire training dataset. Because the subsequent conversation revolves around the drawn goal until it is achieved successfully, it has an important influence on the subsequent training process. Random sampling neglects the way humans acquire knowledge and skills, where they focus on relatively easy materials before harder ones, and hurts the efficiency and stability of the training process. To overcome this problem, Bengio et al. [12] proposed curriculum learning (CL), which presents relatively easy or simple examples at an early stage, inspired by human learning habits. Many NLP tasks have been improved by using a typical CL [13, 14, 15, 16, 17, 18, 19, 20]. These learning methods employed different evaluations of training examples and approaches to adjust the training steps but shared the easy-first strategy. However, in some studies, the reverse version of CL was tested and achieved the best performance among various training scheduler designs [18, 21, 22]. The effectiveness and application of easy-to-hard and hard-to-easy strategies are still worth exploring [23]. Chang et al. [21] stated that the difficult-first strategy is more suitable for cleaner datasets, whereas the classic CL is beneficial for acquiring policies through noisy scenarios and leads to faster convergence. Furthermore, when tasks are difficult for an agent to complete, earlier presentation of easier samples is preferable for an effective training process. However, the exact impact of scheduled training on agent behavior and policy optimization remains unclear. This study aims to improve the performance of a task-oriented dialog system by providing sufficient environment exploration and a training strategy that matches human learning habits. Therefore, we propose Scheduled Curiosity-Deep Dyna-Q (SC-DDQ) which combines CL and a curiosity reward with dialog policy optimization based on Deep Dyna-Q (DDQ), where DDQ is a state-of-the-art model-based dialog system [7], and design a curriculum based on two opposite learning strategies, which are found to be optimal in different scenarios. Compared to count-based exploration approaches [8, 9, 10], and variants of DDQ, SC-DDQ for the first time adapts the curiosity model which models the agent’s intrinsic motivation, and introduces it into DDQ to motivate the agent to explore the environment based on its intrinsic motivation. In addition, SC-DDQ develops criteria for classifying the difficulty of training samples based on the capabilities of the dialog agent and introduces CL into the training of the dialog agent. Because the specific application scenarios of easy-to-difficult and difficult-to-easy learning strategies are not yet conclusive, in the experimental section, we carried out four combinations of experiments based on the presence or absence of the curiosity model and the trend of the difficulty of the task, aiming to explore the effects of agent intrinsic motivation and learning strategies on the completion of the dialog tasks. Preferring more difficult user goals is effective when experiments are conducted without curiosity reward while focusing on easier tasks is found to be an optimal training strategy if the dialog agent selects its action according to both external and internal rewards. To determine and analyze the influences of scheduled training on agent behavior, the entropy of agent action sampling in each training stage, measuring the agent action exploration, was employed to characterize the policy strategy. It is worth noting that entropy is designed to analyze the experimental data, not part of the algorithm itself. We point out that a higher entropy at an earlier stage yields better performance. In other words, at the beginning of learning, relatively uniform action sampling, namely more sufficient action space exploration, leads to a higher task success rate. According to the experimental results, a key factor in facilitating dialog task completion was encouraging action exploration during the early training stage. Compared with [3, 4, 5, 8, 9, 10], SC-DDQ is more scalable. First, this study makes some simplifications to ICM which is the original curiosity model. It performs feature extraction on the state to filter out parts of the state that are unaffected by action execution. Because almost all elements contained in the representation of the dialog state in the context of movie ticket booking scenarios are manipulated by action execution, feature extraction is discarded in this research, which should be borrowed for other scenarios whose states do not contain too many interfering parts. Second, the introduction of CL is mainly an adaptation of the training sequence; therefore, it can be easily extended to algorithms other than dialog systems. Furthermore, our experimental results provide evidence that different training scenarios can potentially benefit from both classical CL and its inverse version. This result presents new ideas for subsequent research on CL. Therefore, the SC-DDQ framework is scalable to real-world applications such as cyber defense. Moradi et al. leveraged reinforcement learning to acquire an attack strategy and allocate defense resources to protect smart electrical power grids[24, 25]. The main idea of SC-DDQ can be adapted to these two cyber defense frameworks. For example, the curiosity reward can be modeled as the prediction error of the defense resource allocation status, and the order of the training tasks presented to the agent can be adjusted according to their defense difficulties. In summary, our main contributions in this paper are three-fold:
@@ 2.1.1 Direct Reinforcement Learning
Dialog policy learning can typically be formulated as a Markov Decision Process (MDP). An episode of task-oriented dialog can be regarded as a set of tuples. In each dialog turn, the policy agent observes the dialog state $s$ and samples action $a$. The selection of action $a$ is based on $\varepsilon$-greed policy, where the action is chosen randomly with probability $\varepsilon$ or to maximize the action-value function $Q(s,a;\theta)$. The action-value function is accomplished by a Multi-Layer Perceptron (MLP) with parameters $\theta_{Q}$. After executing this action, the agent receives a reward $r$ from the environment and observes the user response $a^{u}$. The dialog state is then updated to the next state $s^{\prime}$. The tuple $(s,\ a,\ r,\ a^{u},\ s^{\prime})$ can be viewed as a piece of experience and stored in either a real or a simulated experience replay buffer. The action-value function $Q(\cdot)$ is improved via Deep Q-network (DQN) [6]. An episode of dialog is launched by sampling a user goal from a goal set by the user simulator. Each user goal is defined as G = (inform_slots, request_slots), where inform_slots is a set of constraints and request_slots is a set of requests. For the movie-ticket booking task, typical information slots involve items such as movie names or number of people. Requests can be theater or start time.
Figure 1: The framework of DDQ
@@ 2.2 Curriculum Learning
Curriculum learning (CL) is a training strategy in which the agent initially concentrates on relatively easy data and progressively moves to harder training samples. This strategy is inspired by human study characteristics and is beneficial for accelerating the training convergence speed and enhancing the agent performance. CL is more implementable and compatible than multitudinous and complicated frameworks and models, thereby making it more universally applicable. In addition, CL makes the best use of the existing data, which saves resources and time for RL-based tasks. Typical CL have been used extensively in natural language processing. Liu et al. [17] employed the mastering level of current tasks as the difficulty criterion and guided the agent to grasp easier tasks first. Zhao et al. [19] utilized the total number of inform and request slots for tasks to identify tasks, and the task completion rate to adjust the difficulty level of the next task. However, the easy-to-hard learning strategy is not always helpful. In some studies, the opposite version of CL, focusing on harder examples, brings more significant improvements to the system. Zhao et al. [18] proposed an opposite CL strategy, in which the agent was trained on a general training set and gradually moved to subsets. Moreover, Hacohen et al. [22] demonstrated that self-paced learning, a well-known variant of CL, hurts the agent’s performance. Chang et al. [21] stated that CL is more suitable for training scenarios with significant noise. By contrast, the training process of cleaner scenarios would be more efficient using the opposite CL strategy. As the effectiveness and applicability of difficult-first and easy-first strategies are still open to debate, we utilized the difficulty of user goals as a measurement of dialog complexity and proposed opposing learning schedules based on CL and its reverse version. Applying different reward functions yields a performance boost for the dialog agents based on the two learning schedules.
@@ 2.3 Curiosity Reward
Curiosity is modeled by state error prediction to promote the agent to attempt unconversant states [26, 11]. In RL applications [27, 28, 29, 30, 31], an agent samples an action and receives a reward after observing the current state. Subsequently, the entire dialog environment updates to the next state. If a new state can be predicted accurately before selecting the following action, it is viewed as a familiar state for the agent. However, a state that is difficult to forecast is more informative for the agent. In this study, we integrate the curiosity model with dialog policy optimization to guide agents to explore unfamiliar environment states. In this research, we simplified and adjusted the Intrinsic Curiosity Module (ICM) [11]. ICM is engineered to play video games, such as VizDoom and Super Mario Bros, and it observes screenshots of the game screen as inputs. Because the game screen often contains many disturbances that are not controlled by the action, such as the change of scenery in the background, ICM designs an inverse model to filter out these disturbing factors and thus performs feature extraction on the states. Unlike the application scenarios of ICM, the states in this research are affected by the execution of actions and contain almost no interference factors. Therefore, the curiosity model used in this research removed the inverse model. In addition, ICM calculates the prediction error as the curiosity reward between the next state and the predicted one after the agent executes the action and observes the true next state. In this research, the agent does not execute the feasible actions individually to observe the corresponding next state and calculate the curiosity reward; instead, the agent takes both the curiosity reward and the predicted next state as outputs for training. This eliminates the need for the agent to repeatedly step back, but the disadvantage is that at the early stage of training, the curiosity reward output from our curiosity model may not be accurate enough. Previous studies introduced curiosity-based exploration from a video game scenario into an RL-based dialog system and achieve performance breakthroughs. Wang et al. [32] investigated different exploration strategies, including ICM, in task-completion dialog policy learning and showed improvements. Doering et al. [33] developed a shopkeeper robot whose verbal interactions with customers are guided by curiosity, and it is significantly human-like, compared to non-curious robots. These studies provide evidence for the potential effectiveness of the curiosity model in dialog systems. Therefore, we made some adjustments to the ICM, integrated it into the dialog system, and believe in its effectiveness.
@@ 3.1 Overview
The proposed framework is illustrated in Fig. 3. The SC-DDQ framework consists of three modules: (1) an off-line task classifier for dividing user goals into three complexity levels; (2) a curiosity policy agent for selecting the next action beneficial for completing the task and exploring state by using the current dialog state; and (3) a world model for simulating a user to generate actions and rewards based on real user behaviors. The optimization of SC-DDQ comprises five steps: (1) warm starting: a hand-crafted dialog policy is employed to generate experiences; (2) direct reinforcement learning guided by curiosity: the agent guided by the curiosity model interacts with real users and generates real experiences. They are stored in the real experience replay buffer and used to improve dialog policy; (3) world model learning: the world model is refined on real experiences; (4) planning guided by curiosity: the agent guided by the curiosity model conducts interactions with the world model, where the simulated experiences are collected and employed to optimize dialog policy; and (5) curiosity model training: the curiosity model is optimized using both real and estimated experiences. In addition, the proposed framework can be equipped with different policy models and curriculum schedules. Four combinations are implemented in this research, as shown in Table 1: (1) policy agent without curiosity equipped with the easy-first strategy (EFS) curriculum; (2) policy agent without curiosity equipped with the difficult-first strategy (DFS) curriculum; (3) curiosity-combined policy agent equipped with the EFS curriculum; and (4) curiosity-combined policy agent equipped with the DFS curriculum.
@@
In this research, two opposite types of curriculum training schedules are proposed, both of which enable the RL dialog agent to achieve better performance under certain conditions. To reveal the influence of the training curriculum on the performance of the agent, the response of the agent in each turn of the dialog was recorded. These results are presented in Section 5.2. The structure of the main parts of SC-DDQ is described in the subsequent subsection. The pseudo-code for the iterative SC-DDQ algorithm is shown in Algorithm 1. Line 1 describes the first step in the optimization, where a rule-based dialog is conducted. Lines 2-12 show that the off-line task classifier divides user goals into three levels. Lines 14-17 illustrate that before launching an episode of dialog a user goal that fits the curriculum is sampled. The second step of optimization is described in lines 18-24, where the curiosity policy model outputs Q-values and curiosity rewards and selects the best feasible action based on their sum. Line 24 describes the update of the DQN based on real experiences. The third step of optimization corresponds to line 25 where the world model is updated using real experiences. Line 26 illustrates the fourth step, where the world model, replacing the user simulator, interacts with the agent and generates simulated experiences. Line 27 shows that the DQN is optimized by simulated experiences. The last step of the optimization is shown in line 28, where the curiosity model is updated using both real and simulated experiences.

### v2

@@ 1 Introduction
Since human-computer interaction and natural language processing are in high demand in industry and daily life, and the task-oriented dialog system has become a hot topic and deserves further study and research. Dialog policy model is used to select the best action at each step of a dialog. In a task-oriented dialog system, the goal of the dialog is to complete a specific task, such as booking a movie. The dialog policy should select actions to efficiently achieve the goal.
Dialog policy learning is often formulated as a reinforcement learning (RL) problem [1, 2], in which a dialog agent executes an action based on the observed state and receives a reward from the environment acted by real users. However, optimizing RL agents requires a large number of interactions with the environment, which is an expensive and time-consuming process. Therefore, how to grasp dialog policy efficiently within limited interactions remains a fundamental question for RL [3, 4, 5, 6, 7].
A common approach is to encourage the agent to explore the environment as sufficiently as possible through limited interactions. In contrast to supervised and unsupervised learning, reinforcement learning does not contain a large amount of training data with or without labels prepared in advance, but rather generates experience through the agent’s interactions with the environment and continually optimizes the agent based on feedback from the environment. Therefore, exploration of the environment is crucial for RL, and extensive research has been conducted on promoting agents to explore the environment. Count-based exploration is an intuitive approach in which each state observed by the agent is recorded and the agent is encouraged to learn states with fewer occurrences. The most evident drawback of this approach is its inability to handle complex and high-dimensional data. To address this, Bellemare et al. engineered a model that acquires the distribution of visited environment states to measure the novelty of states and encourage agents to observe unfamiliar states [8]. In addition, to deal with high-dimensional data, Tang et al. suggested mapping high-dimensional states to hash codes [9], and Liu et al. introduced embedding networks to encode the state space [10]. However, a common flaw in such approaches is that they do not lead agents to explore states that have never existed before.
Another method is to model the intrinsic motivation of the agents. One ingenious strategy is to utilize curiosity to facilitate state exploration. Pathak et al. [11] proposed an Intrinsic Curiosity Module (ICM) to model the state prediction error, which depicts the uncertainty and improvement in RL as a curiosity reward to encourage unfamiliar dialog states. This research makes some simplifications and adjustments to ICM to make it more suitable for task-oriented dialog systems to guide unfamiliar dialog state exploration. This model leveraged the predictability of a state to depict its novelty. If the next state can be accurately predicted based on the current state and the upcoming action to be performed, it means that the action cannot lead the agent to see unfamiliar states, and it corresponds to a less curiosity reward. By contrast, if the next state is difficult to predict, the corresponding action receives a larger curiosity reward. The agent executes the action with the largest sum of curiosity rewards and rewards from the environment.
Furthermore, the RL-based dialog agent faces another challenge. Generally, an episode of dialog based on RL is opened up by randomly selecting a user goal from the entire training dataset. Because the subsequent conversation revolves around the drawn goal until it is achieved successfully, it has an important influence on the subsequent training process. Random sampling neglects the way humans acquire knowledge and skills, where they focus on relatively easy materials before harder ones, and hurts the efficiency and stability of the training process. To overcome this problem, Bengio et al. [12] proposed curriculum learning (CL), which presents relatively easy or simple examples at an early stage, inspired by human learning habits. Many NLP tasks have been improved by using a typical CL [13, 14, 15, 16, 17, 18, 19, 20]. These learning methods employed different evaluations of training examples and approaches to adjust the training steps but shared the easy-first strategy.
However, in some studies, the reverse version of CL was tested and achieved the best performance among various training scheduler designs [18, 21, 22]. The effectiveness and application of easy-to-hard and hard-to-easy strategies are still worth exploring [23]. Chang et al. [21] stated that the difficult-first strategy is more suitable for cleaner datasets, whereas the classic CL is beneficial for acquiring policies through noisy scenarios and leads to faster convergence. Furthermore, when tasks are difficult for an agent to complete, earlier presentation of easier samples is preferable for an effective training process. However, the exact impact of scheduled training on agent behavior and policy optimization remains unclear.
This study aims to improve the performance of a task-oriented dialog system by providing sufficient environment exploration and a training strategy that matches human learning habits. Therefore, we propose Scheduled Curiosity-Deep Dyna-Q (SC-DDQ) which combines CL and a curiosity reward with dialog policy optimization based on Deep Dyna-Q (DDQ), where DDQ is a state-of-the-art model-based dialog system [7], and design a curriculum based on two opposite learning strategies, which are found to be optimal in different scenarios. Compared to count-based exploration approaches [8, 9, 10], and variants of DDQ, SC-DDQ for the first time adapts the curiosity model which models the agent’s intrinsic motivation, and introduces it into DDQ to motivate the agent to explore the environment based on its intrinsic motivation. In addition, SC-DDQ develops criteria for classifying the difficulty of training samples based on the capabilities of the dialog agent and introduces CL into the training of the dialog agent.
Because the specific application scenarios of easy-to-difficult and difficult-to-easy learning strategies are not yet conclusive, in the experimental section, we carried out four combinations of experiments based on the presence or absence of the curiosity model and the trend of the difficulty of the task, aiming to explore the effects of agent intrinsic motivation and learning strategies on the completion of the dialog tasks. Preferring more difficult user goals is effective when experiments are conducted without curiosity reward while focusing on easier tasks is found to be an optimal training strategy if the dialog agent selects its action according to both external and internal rewards.
To determine and analyze the influences of scheduled training on agent behavior, the entropy of agent action sampling in each training stage, measuring the agent action exploration, was employed to characterize the policy strategy. It is worth noting that entropy is designed to analyze the experimental data, not part of the algorithm itself. We point out that a higher entropy at an earlier stage yields better performance. In other words, at the beginning of learning, relatively uniform action sampling, namely more sufficient action space exploration, leads to a higher task success rate. According to the experimental results, a key factor in facilitating dialog task completion was encouraging action exploration during the early training stage.
Compared with [3, 4, 5, 8, 9, 10], SC-DDQ is more scalable. First, this study makes some simplifications to ICM which is the original curiosity model. It performs feature extraction on the state to filter out parts of the state that are unaffected by action execution. Because almost all elements contained in the representation of the dialog state in the context of movie ticket booking scenarios are manipulated by action execution, feature extraction is discarded in this research, which should be borrowed for other scenarios whose states do not contain too many interfering parts. Second, the introduction of CL is mainly an adaptation of the training sequence; therefore, it can be easily extended to algorithms other than dialog systems. Furthermore, our experimental results provide evidence that different training scenarios can potentially benefit from both classical CL and its inverse version. This result presents new ideas for subsequent research on CL. Therefore, the SC-DDQ framework is scalable to real-world applications such as cyber defense. Moradi et al. leveraged reinforcement learning to acquire an attack strategy and allocate defense resources to protect smart electrical power grids[24, 25]. The main idea of SC-DDQ can be adapted to these two cyber defense frameworks. For example, the curiosity reward can be modeled as the prediction error of the defense resource allocation status, and the order of the training tasks presented to the agent can be adjusted according to their defense difficulties.
In summary, our main contributions in this paper are three-fold:
@@ 2.1.1 Direct Reinforcement Learning
Dialog policy learning can typically be formulated as a Markov Decision Process (MDP). An episode of task-oriented dialog can be regarded as a set of tuples. In each dialog turn, the policy agent observes the dialog state $s$ and samples action $a$. The selection of action $a$ is based on $\varepsilon$-greed policy, where the action is chosen randomly with probability $\varepsilon$ or to maximize the action-value function $Q(s,a;\theta)$. The action-value function is accomplished by a Multi-Layer Perceptron (MLP) with parameters $\theta_{Q}$. After executing this action, the agent receives a reward $r$ from the environment and observes the user response $a^{u}$. The dialog state is then updated to the next state $s^{\prime}$. The tuple $(s,\ a,\ r,\ a^{u},\ s^{\prime})$ can be viewed as a piece of experience and stored in either a real or a simulated experience replay buffer. The action-value function $Q(\cdot)$ is improved via Deep Q-network (DQN) [6].
An episode of dialog is launched by sampling a user goal from a goal set by the user simulator. Each user goal is defined as G = (inform_slots, request_slots), where inform_slots is a set of constraints and request_slots is a set of requests. For the movie-ticket booking task, typical information slots involve items such as movie names or number of people. Requests can be theater or start time.
@@ 2.2 Curriculum Learning
Curriculum learning (CL) is a training strategy in which the agent initially concentrates on relatively easy data and progressively moves to harder training samples. This strategy is inspired by human study characteristics and is beneficial for accelerating the training convergence speed and enhancing the agent performance. CL is more implementable and compatible than multitudinous and complicated frameworks and models, thereby making it more universally applicable. In addition, CL makes the best use of the existing data, which saves resources and time for RL-based tasks.
Typical CL have been used extensively in natural language processing. Liu et al. [17] employed the mastering level of current tasks as the difficulty criterion and guided the agent to grasp easier tasks first. Zhao et al. [19] utilized the total number of inform and request slots for tasks to identify tasks, and the task completion rate to adjust the difficulty level of the next task.
However, the easy-to-hard learning strategy is not always helpful. In some studies, the opposite version of CL, focusing on harder examples, brings more significant improvements to the system. Zhao et al. [18] proposed an opposite CL strategy, in which the agent was trained on a general training set and gradually moved to subsets. Moreover, Hacohen et al. [22] demonstrated that self-paced learning, a well-known variant of CL, hurts the agent’s performance. Chang et al. [21] stated that CL is more suitable for training scenarios with significant noise. By contrast, the training process of cleaner scenarios would be more efficient using the opposite CL strategy.
As the effectiveness and applicability of difficult-first and easy-first strategies are still open to debate, we utilized the difficulty of user goals as a measurement of dialog complexity and proposed opposing learning schedules based on CL and its reverse version. Applying different reward functions yields a performance boost for the dialog agents based on the two learning schedules.
@@ 2.3 Curiosity Reward
Curiosity is modeled by state error prediction to promote the agent to attempt unconversant states [26, 11]. In RL applications [27, 28, 29, 30, 31], an agent samples an action and receives a reward after observing the current state. Subsequently, the entire dialog environment updates to the next state. If a new state can be predicted accurately before selecting the following action, it is viewed as a familiar state for the agent. However, a state that is difficult to forecast is more informative for the agent. In this study, we integrate the curiosity model with dialog policy optimization to guide agents to explore unfamiliar environment states.
In this research, we simplified and adjusted the Intrinsic Curiosity Module (ICM) [11]. ICM is engineered to play video games, such as VizDoom and Super Mario Bros, and it observes screenshots of the game screen as inputs. Because the game screen often contains many disturbances that are not controlled by the action, such as the change of scenery in the background, ICM designs an inverse model to filter out these disturbing factors and thus performs feature extraction on the states. Unlike the application scenarios of ICM, the states in this research are affected by the execution of actions and contain almost no interference factors. Therefore, the curiosity model used in this research removed the inverse model. In addition, ICM calculates the prediction error as the curiosity reward between the next state and the predicted one after the agent executes the action and observes the true next state. In this research, the agent does not execute the feasible actions individually to observe the corresponding next state and calculate the curiosity reward; instead, the agent takes both the curiosity reward and the predicted next state as outputs for training. This eliminates the need for the agent to repeatedly step back, but the disadvantage is that at the early stage of training, the curiosity reward output from our curiosity model may not be accurate enough.
Previous studies introduced curiosity-based exploration from a video game scenario into an RL-based dialog system and achieve performance breakthroughs. Wang et al. [32] investigated different exploration strategies, including ICM, in task-completion dialog policy learning and showed improvements. Doering et al. [33] developed a shopkeeper robot whose verbal interactions with customers are guided by curiosity, and it is significantly human-like, compared to non-curious robots. These studies provide evidence for the potential effectiveness of the curiosity model in dialog systems. Therefore, we made some adjustments to the ICM, integrated it into the dialog system, and believe in its effectiveness.
@@ 3.1 Overview
The proposed framework is illustrated in Fig. 3. The SC-DDQ framework consists of three modules: (1) an off-line task classifier for dividing user goals into three complexity levels; (2) a curiosity policy agent for selecting the next action beneficial for completing the task and exploring state by using the current dialog state; and (3) a world model for simulating a user to generate actions and rewards based on real user behaviors.
The optimization of SC-DDQ comprises five steps: (1) warm starting: a hand-crafted dialog policy is employed to generate experiences; (2) direct reinforcement learning guided by curiosity: the agent guided by the curiosity model interacts with real users and generates real experiences. They are stored in the real experience replay buffer and used to improve dialog policy; (3) world model learning: the world model is refined on real experiences; (4) planning guided by curiosity: the agent guided by the curiosity model conducts interactions with the world model, where the simulated experiences are collected and employed to optimize dialog policy; and (5) curiosity model training: the curiosity model is optimized using both real and estimated experiences.
In addition, the proposed framework can be equipped with different policy models and curriculum schedules. Four combinations are implemented in this research, as shown in Table 1: (1) policy agent without curiosity equipped with the easy-first strategy (EFS) curriculum; (2) policy agent without curiosity equipped with the difficult-first strategy (DFS) curriculum; (3) curiosity-combined policy agent equipped with the EFS curriculum; and (4) curiosity-combined policy agent equipped with the DFS curriculum.
@@
In this research, two opposite types of curriculum training schedules are proposed, both of which enable the RL dialog agent to achieve better performance under certain conditions. To reveal the influence of the training curriculum on the performance of the agent, the response of the agent in each turn of the dialog was recorded. These results are presented in Section 5.2.
The structure of the main parts of SC-DDQ is described in the subsequent subsection. The pseudo-code for the iterative SC-DDQ algorithm is shown in Algorithm 1. Line 1 describes the first step in the optimization, where a rule-based dialog is conducted. Lines 2-12 show that the off-line task classifier divides user goals into three levels. Lines 14-17 illustrate that before launching an episode of dialog a user goal that fits the curriculum is sampled. The second step of optimization is described in lines 18-24, where the curiosity policy model outputs Q-values and curiosity rewards and selects the best feasible action based on their sum. Line 24 describes the update of the DQN based on real experiences. The third step of optimization corresponds to line 25 where the world model is updated using real experiences. Line 26 illustrates the fourth step, where the world model, replacing the user simulator, interacts with the agent and generates simulated experiences. Line 27 shows that the DQN is optimized by simulated experiences. The last step of the optimization is shown in line 28, where the curiosity model is updated using both real and simulated experiences.